OpenStudy (anonymous):

Solve the differential equation xy+y'=100x

OpenStudy (freckles):

\[y'+xy=100x \\ \text{ integrating factor is } e^{\int_0^x\limits t dt}\] some times people like to call it \[e^{\int\limits x dx}\] but we don't need the constant of integration that is take the constant of integration to be 0

OpenStudy (freckles):

\[v y'+v xy =v 100 x \\ \text{ where we will pretend } v \text{ is what you got for the integrating factor } \\ (vy)'=v100x\] find v replace v then integrate both sides of that equation

OpenStudy (anonymous):

Or \[\frac{y'}{100-y}=x.\]Pick your posion

OpenStudy (freckles):

oh clever didn't even try to see if seperable

OpenStudy (usukidoll):

I was wondering if it was also a separable equation as well

OpenStudy (anonymous):

I'm mainly doing the problem through the separable method.

OpenStudy (usukidoll):

seem like it is xy+y'=100x y'=100x-xy y'=x(100-y) y'/ (100-y) = x and then integrate both sides dy/(100-y) = x dx

OpenStudy (usukidoll):

it's always the rhs side of the equation to be the easiest part

OpenStudy (anonymous):

When I integrate dy/100-y shouldn't I just get ln(100-y)?

OpenStudy (usukidoll):

let's see.. we can use u sub and the log antiderivative